# DetailMaster: Can Your Text-to-Image Model Handle Long Prompts?

### v1

@@ Appendix F Randomness analysis for evaluation with LLMs
To assess the impact of randomness inherent in LLM-based evaluation, we conduct experiments by varying random seeds and setting the LLM temperature to 0.8, subsequently repeating evaluation across five representative models (SD1.5, LongAlign, LLM Blueprint, FLUX.1-dev, and GPT Image-1) to quantify stochastic variability. The results are shown in 9.
@@
The presented results show that the randomness introduced by LLMs in our evaluation framework has minimal impact, with repeated assessments across all models showing negligible variations ($\Delta$ < 0.5%). This robust consistency strongly validates the stability and reliability of our evaluation protocol, ensuring reproducible and dependable model comparisons regardless of inherent LLM randomization factors.
Appendix G Negative correlation between prompt length and attribute alignment
In this section, we analyze the relationship between prompt length and attribute accuracy for detected generated characters (categorized into objects, animals, and persons) to examine how the model’s attribute adherence capability and resistance to attribute misalignment vary with increasing prompt length. As illustrated in Figure 6, the left, middle, and right subgraph respectively show the attribute accuracy trends for “object”, “animal”, and “person” characters.
The trends show a clear inverse correlation between prompt length and attribute accuracy, indicating that longer prompts lead to progressively greater deviations from the intended character descriptions, manifested as attribute misalignment and omission. Notably, the performance comparison reveals that models specifically trained with long prompts (e.g., ELLA and ParaDiffusion) exhibit more gradual accuracy degradation compared to models using conventional training like Deepfloyd, as evidenced by their shallower accuracy decline curves. These findings collectively suggest that current text-to-image generation systems still face substantial challenges in maintaining attribute consistency with long prompts, highlighting the critical need for further research into long-prompt optimization techniques and training methodologies to advance this crucial capability.
Appendix H Mini DetailMaster Benchmark
@@
Appendix I Limitations
@@
Regarding the quality of our benchmark data, we have conducted systematic sampling assessments involving professional annotators, validating the high quality of our dataset, with detailed results presented in Appendix D. Section 5.1 further shows the substantial scale and diversity of our evaluation data, while Table 1 provides comparative analyses highlighting its superior suitability for long-prompt scenarios relative to existing benchmarks. However, our current metrics, while extending T2I-CompBench’s metrics with “Character Locations” and “Scene Attributes”, remain potentially expandable through other dimensions such as aesthetic assessment or human preference scoring. Furthermore, the current prompt collection primarily focuses on realistic imagery and user-uploaded photographs, with limited representation of artistic works. Future work could develop a dedicated subset to evaluate artistic content generation.
@@
Regarding our evaluation protocol, while it leverages LLMs as its core evaluation mechanism, with robustness verified in Appendix F where we demonstrate minimal impact of model randomness on evaluation outcomes, it presents notable computational demands. The current implementation requires substantial GPU resources (minimum 20GB VRAM) and exhibits longer processing times compared to conventional evaluation methods. These resource-intensive characteristics may limit accessibility for some research teams. To address these constraints, we will investigate alternative evaluation methods that can maintain assessment accuracy while improving computational efficiency.
Appendix J Case studies

### v2

@@ Appendix I Robustness of the Single-MLLM Evaluation Protocol
To assess the impact of randomness inherent in LLM-based evaluation, we conduct experiments by varying random seeds and setting the LLM temperature to 0.8, subsequently repeating evaluation across five representative models (SD1.5, LongAlign, LLM Blueprint, FLUX.1-dev, and GPT Image-1) to quantify stochastic variability. The results are shown in 10.
@@
The presented results show that the randomness introduced by LLMs in our evaluation framework has minimal impact, with repeated assessments across all models showing negligible variations ($\Delta$ << 0.5%). This robust consistency strongly validates the stability and reliability of our evaluation protocol, ensuring reproducible and dependable model comparisons regardless of inherent LLM randomization factors.
Appendix K Compatibility of Our Framework with Other Evaluation Metrics
Our work specifically focuses on the adherence of T2I models to detailed, long prompts. Consequently, metrics such as “visual quality” and “safety” fall outside the scope of our objectives. Nevertheless, our benchmark is fully compatible with other frameworks that assess “visual quality” and “safety” scores. The prompts and generated images from our dataset can be readily integrated into such frameworks to obtain these corresponding scores.
To demonstrate this compatibility, we evaluate our prompts and generated images using several established external metrics: (1) DiffSynth-Studio’s ImageReward (for image quality), Aesthetic (for aesthetic scores), and HPSv2.1 and MPS (for human preference scores) (Xu et al., 2023; Duan et al., 2024; Wu et al., 2023b; Zhang et al., 2024); and (2) Falcons-ai’s nsfw_image_detection (for safety assessment) (Falcons.ai, 2023).
Table 11: Demonstration of compatibility with external evaluation frameworks.
Model ImageReward Aesthetic HPSv2.1 MPS nsfw_image_detection
SD-XL 0.08 5.67 0.2831 9.69 0.0011
FLUX.1-dev 0.45 5.41 0.2921 9.89 0.0005
ELLA 0.23 5.52 0.2486 9.18 0.0021
ParaDiffusion 0.16 5.60 0.2733 8.48 0.0011
As shown in Table 11, our benchmark can be effectively combined with other evaluation frameworks to obtain other metrics. In our code repository, we will include hyperlinks to these external evaluation frameworks, providing users with the flexibility to choose additional metrics according to their specific needs.
Furthermore, the method’s ability to extract fine-grained information from individual image entities makes it a powerful tool for constructing datasets and benchmarks for guided data-model co-development (Qin et al., 2025). Applications include discerning subtle differences between similar images (Jiao et al., 2025), capturing nuanced human attributes like actions and expressions (Zhou et al., 2024), generating Visual Question Answering (VQA) samples (Kuang et al., 2025b), and annotating unconventional data types, such as images containing emojis or web GUIs (Kuang et al., 2025a; Luo et al., 2025).
Appendix L Mini DetailMaster Benchmark
@@
Appendix M Statistical Analysis of Metric Difficulty
In this section, we present an experimental analysis of the relative difficulty of our four metrics. Furthermore, we conduct a failure case analysis to investigate instances where the evaluated models perform inaccurately.
Specifically, for each sample, we calculate the proportion of correctly generated attributes for each evaluation metric. If a model achieves a correct proportion below 50% for a specific metric, we consider that metric hard for the model on that sample (e.g., if a sample’s prompt contains eight “animal attributes” and the model generates fewer than four correct “animal attributes”, we consider the “Animal” generation task hard for the model on that sample).
The evaluation metrics we examine include: “Object”, “Animal”, “Person”, “Character Locations”, “Scene Attributes”, and “Entity Relationships”. Notably, the “Character Attributes” metric in our main evaluation is split into “Object”, “Animal”, and “Person” for individual analysis. Meanwhile, the sub-metrics of “Scene Attributes” (i.e., “Background”, “Light”, and “Style”) are aggregated.
The results present the proportion of hard samples for each metric. We also calculate the average value across all evaluated models, as presented in the “Average” row. The final outcomes are shown in Table 13.
Table 13: Proportion of hard samples for each metric.
Model Character Attributes Character Locations Scene Attributes Entity Relationships
Object Animal Person
SD1.5 82.6% 75.7% 88.2% 95.8% 38.6% 98.1%
SD-XL 79.0% 74.0% 85.2% 94.2% 40.8% 95.7%
DeepFloyd IF 72.1% 64.8% 77.1% 92.3% 33.3% 94.6%
SD3.5 Large 54.4% 56.0% 69.6% 76.5% 2.7% 62.9%
FLUX.1-dev 51.7% 56.5% 67.9% 68.8% 0.8% 52.2%
Gemini 2.0 Flash 47.1% 56.2% 68.9% 65.8% 0.6% 47.4%
GPT Image-1 42.7% 55.6% 62.9% 55.2% 0.2% 32.4%
LLM4GEN 81.7% 73.8% 84.9% 95.3% 52.8% 97.3%
LLM Blueprint 82.0% 74.5% 87.8% 88.3% 29.4% 94.5%
ELLA 69.0% 68.5% 79.9% 90.9% 43.6% 90.4%
LongAlign 71.2% 68.1% 86.4% 91.9% 11.6% 87.0%
ParaDiffusion 68.3% 68.6% 80.1% 88.1% 9.8% 80.0%
Average 66.8% 66.0% 78.2% 83.6% 22.0% 77.7%
As shown in the table, the most challenging metric is “Character Locations”. We check cases where all models made errors on this metric and identify two main reasons for the failures: (1) The object positions described in the prompt may contradict conventional real-world positions in photos. For instance, when the prompt states that traffic lights are in the “lower part” of the image while they typically appear in the “upper part” in real photos, this discrepancy leads to positional misalignment. (2) The models exhibit a tendency to generate objects in the middle of the image, even when the prompt specifies left or right placement.
The second most challenging metric is “Person” under the “Character Attributes” metric. From the error cases, we observe two failure patterns: (1) descriptions of the person often involve interactions with other objects (e.g., “holding a cat”); and (2) the clothing descriptions for the person are complex.
Regarding “Entity Relationships”, the error cases are often correlated with the “Character Attributes” issues. When an object’s attributes contain excessive errors, the evaluator will not recognize the object’s existence, leading to negative judgments. Additionally, errors occur when the object’s position deviates from conventional real-world placements.
Regarding “Scene Attributes”, this metric is relatively the simplest. Models typically generate accurate results as long as the supported prompt length is sufficient. For erroneous cases, aside from prompt truncation issues, failures primarily occur when: (1) the background contains excessive structural details, or (2) the style descriptions are specialized or uncommon.
Appendix N Negative Correlation between Prompt Length and Attribute Alignment
In this section, we analyze the relationship between prompt length and attribute accuracy for the detected generated characters (categorized into objects, animals, and persons) to examine how the model’s attribute adherence vary with increasing prompt length. As illustrated in Figure 6, the left, middle, and right subgraph respectively show the attribute accuracy trends for “object”, “animal”, and “person” characters.
The trends show a clear inverse correlation between prompt length and attribute accuracy, indicating that longer prompts lead to progressively greater deviations from the intended character descriptions, manifested as attribute misalignment and omission. Notably, the performance comparison reveals that models specifically trained with long prompts (e.g., ELLA and ParaDiffusion) exhibit more gradual accuracy degradation compared to models using conventional training like Deepfloyd, as evidenced by their shallower accuracy decline curves. These findings collectively suggest that current text-to-image generation still face substantial challenges in maintaining attribute consistency under detail-intensive conditions, highlighting the critical need for further research into long-prompt optimization techniques and training methodologies to advance this crucial capability.
Appendix O Limitations
@@
Regarding the quality of our benchmark data, we have conducted systematic sampling assessments involving professional annotators, validating the high quality of our dataset (detailed results presented in Appendix G). Section 5.1 further shows the substantial scale and diversity of our data, and Table 1 also provides comparative analyses highlighting its superior suitability for long-prompt scenarios relative to existing benchmarks. However, our current metrics, while extending T2I-CompBench’s metrics with “Character Locations” and “Scene Attributes”, remain potentially expandable through other dimensions such as aesthetic assessment or human preference scoring.
@@
Regarding our evaluation protocol, it leverages LLMs as its core evaluation mechanism, which presents notable computational demands. The current implementation requires substantial GPU resources (minimum 20GB VRAM) and exhibits longer processing times compared to conventional evaluation methods. These resource-intensive characteristics may limit accessibility for some research teams. To address these constraints, we will investigate alternative evaluation methods that can maintain assessment accuracy while improving computational efficiency.
Appendix P Case Studies